# Mind the Gap Why Neural Memory Fails Under Semantic Density

### v1

@@ Abstract
The brain solves a problem that current AI architectures do not: storing specific episodic facts without corrupting general semantic knowledge. Neuroscience explains this capacity through Complementary Learning Systems theory, which posits that biological memory requires two distinct subsystems - a fast hippocampal system for episodic storage using sparse, pattern-separated representations, and a slow neocortical system for extracting statistical regularities through distributed, overlapping representations. Current AI systems lack this architectural separation, attempting to serve both functions through neural weights alone, and we argue that this approach faces fundamental limits.
We identify what we term the Stability Gap in online neural memory: fast-weight and associative mechanisms that write facts into shared continuous parameters at inference time collapse to near-random accuracy within tens of semantically related facts. Through the lens of semantic density ($\rho$), we show that this collapse occurs as few as $N=5$ facts at high density ($\rho>0.6$) or $N\approx 20$7575 at moderate density, a phenomenon we formalize as the Orthogonality Constraint. Critically, this failure is not about attention capacity or context length - it persists even with perfect attention and unlimited context, arising instead from write-time interference when storage and retrieval share the same continuous substrate.
[…]
We propose Knowledge Objects (KOs) as the missing hippocampal component: discrete, typed memory units with controlled vocabularies and explicit version chains that provide interference-free episodic storage. Paired with neural weights serving the neocortical role, Knowledge Objects enable a true complementary learning architecture, and our findings suggest that reliable AI memory may require this bicameral design.
@@
neural memory, associative memory, complementary learning systems, episodic memory, retrieval-augmented generation, knowledge representation, transformers
@@ 1 Introduction
The brain maintains various distinct memory systems, including the hippocampus (which rapidly encodes specific episodes with minimal interference), and the neocortex (which can be thought to slowly extract statistical patterns across many experiences). This division is not an accident of biology, but a solution to a fundamental computational problem; the same representations cannot simultaneously support fast, precise episodic storage and slow, distributed semantic learning (McClelland et al., 1995). Current AI architectures lack this separation, attempting to handle both episodic facts and semantic knowledge within the same neural substrate - whether context windows, parametric weights, or fast-weight associative memory. We argue that for online episodic storage via inner-product-addressed superposition in shared continuous parameters, overlap-driven interference imposes fundamental limits unless the system introduces explicit pattern separation.
@@
This paper distinguishes four forms of “memory” in AI systems that are often conflated in the literature. The first is context-window attention, where facts remain visible within the current input sequence; this form is volatile (lost when the session ends) and incurs cost that grows with the amount of remembered material. The second is slow parametric memory in model weights, where facts are compressed into parameters during pretraining or finetuning; this form offers vast capacity but is expensive to update and does not natively support surgical overwrite, deletion, or version semantics. The third is fast parametric memory via online or associative mechanisms, where new facts are written at inference time into shared continuous parameters; this form enables instant writes but proves susceptible to interference under semantic density. The fourth is discrete addressable storage, where facts are stored as explicit key-value records in external data structures; this form supports deterministic updates and avoids interference between entries entirely. Our analysis demonstrates that the first three forms cannot reliably support precise episodic memory under realistic conditions, and that the fourth—discrete addressing—is architecturally necessary for the hippocampal function.
The Stability Gap.
@@
We identify a core failure mode that we term the Stability Gap: online neural memory collapses under semantic density. When facts share semantic overlap—for instance, “Paris is the capital of France” and “Lyon is a city in France”—their representations interfere during retrieval because their embedding vectors occupy nearby regions of the continuous space. We formalize this phenomenon as the Orthogonality Constraint and demonstrate empirically that fast-weight and associative mechanisms collapse to near-random accuracy within tens of semantically related facts, far below theoretical capacity bounds that assume orthogonal keys. This limitation is not a bug to be fixed with scale or architectural refinement; it is a geometric inevitability arising from inner-product retrieval in shared continuous storage, where the training objective actively encourages the semantic clustering that causes storage interference.
Production Failures.
Production systems have partially recognized this problem by adopting discrete external databases for cross-session memory—when users enable “memory” features in ChatGPT or Claude, their facts are stored in databases rather than encoded into model weights. However, current implementations suffer from two failure modes that explain why memory can appear unreliable even when retrieval accuracy appears adequate. Schema drift occurs when predicate naming is inconsistent across extractions, with our experiments showing only 40–70% schema match across models. Version ambiguity occurs when corrections fail to cleanly supersede prior values, with some models mentioning both old and new values rather than providing the canonical current state.
@@
The Orthogonality Constraint applies to any system that attempts online storage of discrete facts in shared continuous parameters. It does not claim that LLMs cannot store facts; they obviously can, via slow pretraining. The constraint specifically concerns inference-time episodic memory. We do not claim that attention mechanisms are unnecessary, nor that larger models cannot memorize more facts during pretraining. Rather, we show that attention alone does not provide stable write-time memory under semantic interference—no regime exists in which increased attention capacity alone improves long-term factual retention under semantic density. Similarly, we do not claim that RAG is useless; we show that naive similarity-only RAG lacks the identity, schema, and version semantics required for true episodic memory, and that adding these capabilities effectively introduces KO-like structure.
@@
Figure 1: The Memory Architecture Gap. (a) Pure neural approaches to online memory—context windows and fast weights—face the Orthogonality Constraint: interference grows with semantic density, causing collapse. Production systems have adopted discrete external storage (databases) but suffer from schema drift and version ambiguity. (b) Our hybrid architecture provides structured discrete storage (Knowledge Objects) with controlled vocabularies and version chains, connected to the LLM via a learned router. This explains why discrete storage is necessary and fixes the remaining failure modes.
@@ 2 Related Work
Stability Gap in Continual Learning.
The term “Stability Gap” has been used in continual learning to describe the drop in general capabilities when fine-tuning on new domains (Guo et al., 2024). Our usage is distinct: we define the Semantic Stability Gap as the precise threshold ($N$) at which episodic retrieval collapses to near-random accuracy due to interference in embedding space. The continual learning stability gap concerns capability degradation (forgetting how to perform general tasks); our stability gap concerns fact collision (inability to retrieve specific facts when semantically similar facts crowd the same vector space region).
@@
Recent architectures like Titans (Munkhdalai et al., 2025) and Linear Transformers (Katharopoulos et al., 2020) learn “fast weights” at test time, treating memory as gradient optimization. While Titans reports 98.8% accuracy on associative recall, it provides no capacity bounds and does not characterize failure modes under semantic density. We show that the core mechanism underlying such architectures—superposition of key-value pairs in shared continuous parameters—leads to collapse when storing semantically related facts.
@@
Our architecture is directly inspired by Complementary Learning Systems (CLS) theory from neuroscience (McClelland et al., 1995; Kumaran et al., 2016). CLS proposes that biological memory requires two systems with fundamentally different properties: a hippocampal system characterized by fast learning, sparse and discrete representations, and pattern separation that enables storage of specific episodes without interference; and a neocortical system characterized by slow learning, distributed and overlapping representations, and pattern completion that extracts statistical regularities through gradual exposure across many experiences. The hippocampus solves exactly the problem we address: storing arbitrary new facts quickly without corrupting existing memories, which it achieves through sparse, non-overlapping representations—the biological analog of discrete addressing. Our Knowledge Objects play the hippocampal role, while neural weights play the neocortical role, and sleep consolidation (which transfers hippocampal memories to cortex) maps to our proposed mechanism for promoting stable KO facts into model weights.
@@ 3 Theoretical Framework: The Orthogonality Constraint
The problem is that semantic embeddings cannot exploit this geometric capacity. The training objective for embedding models requires similar concepts to cluster together—this is what enables generalization and transfer. “Paris is the capital of France” and “Lyon is a city in France” must be nearby in embedding space because they share words, appear in similar contexts, and predict similar continuations.
The Orthogonality Constraint thus arises not from geometric limits on $d$, but from the fundamental tension between two objectives that cannot be simultaneously satisfied. The compression objective—what embeddings are trained for—requires grouping semantically similar content together to enable generalization and transfer learning. The storage objective—what factual recall requires—demands separating distinct facts to enable interference-free retrieval. These objectives are fundamentally opposed: the same geometric property (clustering similar concepts) that makes embeddings useful for prediction makes them unreliable for storage.
@@
The embedding model is trained to put “Paris is capital of France” near “Lyon is city in France” because they share words, appear in similar contexts, and predict similar continuations. This clustering is a feature for generalization—it enables the model to transfer knowledge about France to new contexts. But it is a bug for factual lookup—it causes “Paris” and “Lyon” to interfere with each other during retrieval.
If pretraining had successfully orthogonalized facts, we would not observe: (1) the “Lost in the Middle” phenomenon, (2) our measured stability gap collapse, (3) models hallucinating similar-but-wrong facts, or (4) semantic drift in long-context agents. But we observe all of these. The JL capacity exists in theory, but the training objective does not exploit it for fact storage—it exploits it for pattern compression.
Real-world knowledge exhibits $\rho\approx 0.3$$0.7$, not $\rho\approx 0$. This is why collapse occurs at $N\approx 20$—far below both the JL bound and the classical Hopfield capacity of $\sim 0.14d\approx 54$ for $d=384$.
@@
In discrete memory, keys are hash addresses; if $\mathbf{k}\neq\mathbf{k}_{\text{old}}$, overlap is exactly 0 by construction. The Stability Gap—the accuracy difference between discrete and neural memory—is thus a direct consequence of accumulated interference in continuous representations.
Figure 2: The Orthogonality Constraint. Left: In neural memory, semantically related facts (here, two statements about French cities) produce embedding vectors with small angular separation ($\cos\theta\approx 0.7$). When retrieving one fact, the system also partially retrieves others with similar keys—interference that grows with corpus size $N$ and semantic density $\rho$, causing collapse by $N\approx 20$7575 depending on density. Right: Discrete memory assigns unique hash addresses to each fact; overlap between addresses is exactly zero by construction. Retrieval returns $\mathbf{v}_{A}$ with no interference regardless of how many other facts share semantic similarity.
@@ 4.1 Neural Memory Baseline: Linear Associative Memory
where $\lambda=0.99$ is a decay factor and $\eta=0.1$ is the learning rate. Retrieval proceeds by computing $\hat{\mathbf{v}}=\mathbf{M}\mathbf{q}$ for a query $\mathbf{q}$ and decoding via nearest-neighbor matching against stored values. This baseline captures the core failure mode we analyze: gradient updates to shared weights cause interference when keys overlap in the embedding space. Production systems may incorporate additional mechanisms such as normalization and regularization, but the fundamental constraint—that semantically similar keys produce blended retrievals—remains inherent to any architecture that stores facts via superposition.

### v2

# Attention Is Not Retention: The Orthogonality Constraint in Infinite-Context Architectures

@@ Abstract
The mammalian brain has developed a solution to a problem that current AI architectures are yet to solve: storing specific episodic facts without corrupting general semantic knowledge. Neuroscience explains this capacity through Complementary Learning Systems theory, which posits that biological memory requires two distinct subsystems, a fast hippocampal system for episodic storage using sparse, pattern-separated representations, and a slow neocortical system for extracting statistical regularities through distributed, overlapping representations. Current AI systems lack this architectural separation, attempting to serve both functions through neural weights alone.
We identify a fundamental geometric limit we term the Orthogonality Constraint: reliable memory requires orthogonal keys, but semantic embeddings cannot be orthogonal because training clusters similar concepts together. The result is Semantic Interference (connecting to what cognitive psychologists have long observed in human memory retrieval), where online neural memory systems writing facts into shared continuous parameters experience interference proportional to semantic density, collapsing to near-random accuracy within tens of semantically related facts. Though we use “semantic” to link to this literature, the interference mechanism operates in any embedding space where training clusters similar items. Through the lens of semantic density ($\rho$), the mean pairwise cosine similarity of embeddings, we show that interference causes collapse at as few as $N=5$ facts at high density ($\rho>0.6$) or $N\approx 20$7575 at moderate density. We validate this constraint across domains and modalities: 16,309 Wikipedia facts (real-world text), scientific measurements ($\rho=0.96$, neural accuracy 0.02% at N=10, [^0] N=10{,}000), and image embeddings ($\rho_{\text{intra}}=0.82$, neural accuracy 0.05% at N=2, [^0] N=2{,}000). This failure is geometric, not architectural; no regime exists in which increased model capacity, attention span, or training data can overcome interference when keys share semantic overlap.
The Orthogonality Constraint has immediate architectural consequences: reliable episodic memory requires discrete addressing. We propose Knowledge Objects (KOs), structured facts with hash-based identity, controlled vocabularies, and explicit version chains, as the architecture that satisfies this constraint. On 16,309 Wikipedia facts, KO embedding retrieval achieves 45.7% accuracy where Modern Hopfield Networks (the mathematical foundation of transformer attention) collapse to near-zero; hash-based retrieval maintains 100% throughout. Production systems (Claude Memory, ChatGPT Memory) have partially adopted discrete storage but store unstructured text, causing schema drift (40–70% predicate consistency) and version ambiguity (0–100% clean correction rates). Knowledge Objects address these limitations, providing the complete hippocampal component that current systems lack. Paired with neural weights for semantic reasoning, this bicameral architecture resolves the tension between fast episodic storage and slow statistical learning that pure neural approaches cannot escape.
@@
neural memory, associative memory, complementary learning systems, episodic memory, retrieval-augmented generation, knowledge representation, neurosymbolic AI, transformers
@@ 1 Introduction
The brain separates two functions that AI architectures conflate: a hippocampus that rapidly encodes specific episodes with minimal interference, and a neocortex that slowly extracts statistical patterns across many experiences. This division is not an accident of biology but a solution to a fundamental computational problem, the same representations cannot simultaneously support fast, precise episodic storage and slow, distributed semantic learning (McClelland et al., 1995). Current AI architectures lack this separation, attempting to handle both episodic facts and semantic knowledge within the same neural substrate, whether context windows, parametric weights, or fast-weight associative memory. We argue that for online episodic storage via inner-product-addressed superposition in shared continuous parameters, overlap-driven interference imposes fundamental limits unless the system introduces explicit pattern separation.
@@
This paper distinguishes four forms of “memory” in AI systems that are often conflated in the literature. The first is context-window attention, where facts remain visible within the current input sequence; this form is volatile (lost when the session ends) and incurs cost that grows with the amount of remembered material. The second is slow parametric memory in model weights, where facts are compressed into parameters during pretraining or finetuning; this form offers vast capacity but is expensive to update and does not natively support surgical overwrite, deletion, or version semantics. The third is fast parametric memory via online or associative mechanisms, where new facts are written at inference time into shared continuous parameters; this form enables instant writes but proves susceptible to interference under semantic density. The fourth is discrete addressable storage, where facts are stored as explicit key-value records in external data structures; this form supports deterministic updates and avoids interference between entries entirely. Our analysis demonstrates that the first three forms cannot reliably support precise episodic memory under realistic conditions, and that the fourth, discrete addressing, is architecturally necessary for the hippocampal function.
Semantic Interference.
@@
We identify a core failure mode we term Semantic Interference: online neural memory collapses under semantic density. When facts share semantic overlap, for instance, “Paris is the capital of France” and “Lyon is a city in France”, their representations interfere during retrieval because their embedding vectors occupy nearby regions of the continuous space. We formalize this phenomenon as the Orthogonality Constraint and demonstrate empirically that fast-weight and associative mechanisms collapse to near-random accuracy within tens of semantically related facts, far below theoretical capacity bounds that assume orthogonal keys. This limitation is not a bug to be fixed with scale or architectural refinement; it is a geometric inevitability arising from inner-product retrieval in shared continuous storage, where the training objective actively encourages the semantic clustering that causes storage interference.
Production Systems: Partial Solutions.
Production systems have recognized this problem and moved toward discrete storage. ChatGPT Memory stores short timestamped entries like “User is allergic to shellfish” [^1] OpenAI Help Center, “Memory FAQ,” https://help.openai.com/en/articles/8590148-memory-faq while Claude Memory stores memories in Markdown files. [^2] Anthropic documentation describes a file-based architecture using CLAUDE.md files; see https://docs.anthropic.com and community analyses. This represents progress, but only partial progress. These systems store unstructured text extracted by LLMs rather than structured facts, and this causes two critical failure modes that we document empirically. First, schema drift: predicate naming is inconsistent across extractions, the same fact might be stored as “likes Python,” “prefers Python,” or “favorite language is Python” depending on phrasing, breaking downstream lookup. Second, version ambiguity: corrections fail to cleanly supersede prior values, with some models mentioning both old and new values rather than providing the canonical current state. Our experiments show 40–70% schema consistency and 0–100% clean correction rates across frontier models. Production systems have escaped the Orthogonality Constraint by adopting discrete storage, but they have not yet built the structured semantics required for reliable operation at scale.
@@
The Orthogonality Constraint applies to any system that attempts online storage of discrete facts in shared continuous parameters. It does not claim that LLMs cannot store facts; they obviously can, via slow pretraining. The constraint specifically concerns inference-time episodic memory. We do not claim that attention mechanisms are unnecessary, nor that larger models cannot memorize more facts during pretraining. Rather, we show that attention alone does not provide stable write-time memory under semantic interference; no regime exists in which increased attention capacity alone improves long-term factual retention under semantic density. Similarly, we do not claim that RAG is useless; we show that naive similarity-only RAG lacks the identity, schema, and version semantics required for true episodic memory, and that adding these capabilities effectively introduces KO-like structure.
@@
Figure 1: The Memory Architecture Gap. (a) Pure neural approaches to online memory, context windows and fast weights, face the Orthogonality Constraint: interference grows with semantic density, causing collapse. Production systems have adopted discrete external storage but use unstructured text, causing schema drift and version ambiguity. (b) Knowledge Objects provide structured discrete storage with hash-based identity, controlled vocabularies, and version chains. This is the architecture that satisfies the Orthogonality Constraint while providing the semantic guarantees that production systems lack.
@@ 2 Related Work
Semantic Interference in Human and Machine Memory.
We term our core phenomenon Semantic Interference, connecting to what cognitive psychologists have long observed in human memory retrieval: semantically similar items interfere with each other during recall (Loewenstein et al., 2004; Müller & Pilzecker, 1900). Müller and Pilzecker identified this effect in 1900; over a century of subsequent research has confirmed that retrieval accuracy degrades when target items share semantic features with competitors. Our contribution is providing the mathematical framework explaining why this occurs in any associative memory system: the orthogonality constraint (Eq. 2) is violated when semantically similar items cluster in embedding space.
@@
Recent architectures like Titans (Munkhdalai et al., 2025) and Linear Transformers (Katharopoulos et al., 2020) learn “fast weights” at test time, treating memory as gradient optimization. While Titans reports 98.8% accuracy on associative recall, it provides no capacity bounds and does not characterize failure modes under semantic density. We show that the core mechanism underlying such architectures, superposition of key-value pairs in shared continuous parameters, leads to collapse when storing semantically related facts.
Nested Learning and Associative Memory.
Behrouz et al. (Behrouz et al., 2025) formalize the view that neural architectures and optimizers are nested systems of associative memories, each compressing its own “context flow” at different update frequencies. Their framework elegantly unifies learning dynamics across multiple timescales, drawing on the same neuroscientific inspiration (brain oscillations, multi-timescale processing) that motivates our work. However, their analysis assumes that effective memory can be achieved through learned compression in continuous parameters. Our Orthogonality Constraint identifies a fundamental limit they do not address: when stored keys share semantic similarity, interference during retrieval becomes geometrically inevitable regardless of how the memory is trained. Their “Continuum Memory System” proposes a spectrum of update frequencies; our contribution is explaining why at least one level of this spectrum must employ discrete addressing to achieve interference-free episodic storage.
@@
Our architecture is directly inspired by Complementary Learning Systems (CLS) theory from neuroscience (McClelland et al., 1995; Kumaran et al., 2016). CLS proposes that biological memory requires two systems with fundamentally different properties: a hippocampal system characterized by fast learning, sparse and discrete representations, and pattern separation that enables storage of specific episodes without interference; and a neocortical system characterized by slow learning, distributed and overlapping representations, and pattern completion that extracts statistical regularities through gradual exposure across many experiences. The hippocampus solves exactly the problem we address: storing arbitrary new facts quickly without corrupting existing memories.
Critically, the hippocampus achieves this through a specific mechanism: the Dentate Gyrus performs pattern separation, taking overlapping sensory inputs and mapping them to sparse, orthogonal neural representations. This forces semantically similar inputs (“kitchen today” vs. “kitchen yesterday”) onto distinct neural populations, preventing interference. Our hash function $\text{hash}(\text{subject}\|\text{predicate})$ serves exactly this role: it takes semantically dense embedding vectors (which overlap in continuous space) and maps them to discrete addresses (which are orthogonal by construction). Knowledge Objects thus implement an artificial Dentate Gyrus, a pattern separation layer that overcomes the Orthogonality Constraint by forcing storage into non-interfering addresses. Neural weights play the neocortical role of slow semantic generalization, and sleep consolidation (which transfers hippocampal memories to cortex) maps to our proposed mechanism for promoting stable KO facts into model weights.
@@ 3 Theoretical Framework: The Orthogonality Constraint
Information-Theoretic Interpretation.
Eq. (2) admits a signal-to-noise interpretation: the target fact contributes signal $\mathbf{v}_{j}(\mathbf{k}_{j}\cdot\mathbf{k}_{j})$, while all other facts contribute interference noise $\sum_{i\neq j}\mathbf{v}_{i}(\mathbf{k}_{i}\cdot\mathbf{k}_{j})$. The retrieval signal-to-noise ratio (SNR) thus scales inversely with $N\cdot\rho^{2}$. This connects Semantic Interference to Shannon-Hartley channel capacity: just as a communication channel fails when noise power exceeds signal power, associative memory fails when interference overwhelms the target. The critical difference from classical Hopfield capacity bounds is that we measure semantic density rather than assuming random patterns: semantic embeddings, by design, exhibit far higher density than random vectors would.
@@
The problem is that semantic embeddings cannot exploit this geometric capacity. The training objective for embedding models requires similar concepts to cluster together, this is what enables generalization and transfer. “Paris is the capital of France” and “Lyon is a city in France” must be nearby in embedding space because they share words, appear in similar contexts, and predict similar continuations.
The Orthogonality Constraint thus arises not from geometric limits on $d$, but from the fundamental tension between two objectives that cannot be simultaneously satisfied. The compression objective, what embeddings are trained for, requires grouping semantically similar content together to enable generalization and transfer learning. The storage objective, what factual recall requires, demands separating distinct facts to enable interference-free retrieval. These objectives are fundamentally opposed: the same geometric property (clustering similar concepts) that makes embeddings useful for prediction makes them unreliable for storage.
@@
The embedding model is trained to put “Paris is capital of France” near “Lyon is city in France” because they share words, appear in similar contexts, and predict similar continuations. This clustering is a feature for generalization, it enables the model to transfer knowledge about France to new contexts. But it is a bug for factual lookup, it causes “Paris” and “Lyon” to interfere with each other during retrieval.
If pretraining had successfully orthogonalized facts, we would not observe: (1) the “Lost in the Middle” phenomenon, (2) the interference collapse we document, (3) models hallucinating similar-but-wrong facts, or (4) semantic drift in long-context agents. But we observe all of these. The JL capacity exists in theory, but the training objective does not exploit it for fact storage, it exploits it for pattern compression.
Real-world knowledge exhibits $\rho\approx 0.3$$0.7$, not $\rho\approx 0$. This is why collapse occurs at $N\approx 20$, far below both the JL bound and the classical Hopfield capacity of $\sim 0.14d\approx 54$ for $d=384$.
@@
In discrete memory, keys are hash addresses; if $\mathbf{k}\neq\mathbf{k}_{\text{old}}$, overlap is exactly 0 by construction. Semantic Interference (the accuracy gap between discrete and neural memory) is thus a direct consequence of accumulated interference in continuous representations.
Figure 2: The Orthogonality Constraint. Left: In neural memory, semantically related facts (here, two statements about French cities) produce embedding vectors with small angular separation ($\cos\theta\approx 0.7$). When retrieving one fact, the system also partially retrieves others with similar keys, interference that grows with corpus size $N$ and semantic density $\rho$, causing collapse by $N\approx 20$7575 depending on density. Right: Discrete memory assigns unique hash addresses to each fact; overlap between addresses is exactly zero by construction. Retrieval returns $\mathbf{v}_{A}$ with no interference regardless of how many other facts share semantic similarity.
@@ 4.1 Neural Memory Baseline: Linear Associative Memory
where $\lambda=0.99$ is a decay factor and $\eta=0.1$ is the learning rate. Retrieval proceeds by computing $\hat{\mathbf{v}}=\mathbf{M}\mathbf{q}$ for a query $\mathbf{q}$ and decoding via nearest-neighbor matching against stored values. This baseline captures the core failure mode we analyze: gradient updates to shared weights cause interference when keys overlap in the embedding space. Production systems may incorporate additional mechanisms such as normalization and regularization, but the fundamental constraint, that semantically similar keys produce blended retrievals, remains inherent to any architecture that stores facts via superposition.